# Inference in Tightly Identified and Large-Scale Sign-Restricted SVARs

### 24 Apr 2026 v1

@@ 5.1 Small Oil Market Model
As an example of a large SVAR model, we consider a 35-variable quarterly model of the U.S. economy with five lags and a constant. This application extends the work of [chan2025], which in turn builds on [crump2021]. The sample period spans from 1975:Q2 to 2020:Q4.111111We thank Christian Matthes for sharing the replication files. The variables are listed in Table 2, which also gives the identifying restrictions. Following [arias2025elliptical], both sign and ranking restrictions are imposed to identify ten structural shocks: demand, investment, financial, monetary, government spending, technology, labor supply, wage bargaining, oil price, and consumer sentiment shocks.
Following [arias2025elliptical], we place a Minnesota prior on the reduced-form parameters. [^12] We thank Jonas Arias for sharing their prior with us. As the model contains only contemporaneous restrictions, we estimate it using the non-centered specification, where instead of sampling the VAR coefficient matrices, we sample an auxiliary standard normal vector and transform it into the desired parameters $\tilde{A}$. Further details are provided in the Appendix D.
@@
Figure 8 depicts the impulse response functions of the real GDP, Fed funds rate, unemployment rate, PCE index, nonresidential investment, and real wage to a one standard deviation demand and investment shocks. The impulse responses are visually identical to those obtained using the elliptical slice sampler of [arias2025elliptical] under the same prior distribution and identifying restrictions (as in the previous section, we implemented their sampler in R, but report times based on a faster Matlab implementation). This suggests that our sampling algorithm correctly targets the same posterior distribution. In particular, the left panel of Figure 8 shows that the demand shock causes a temporary increase in the real GDP, prices, and the federal funds rate, but reduces unemployment. The demand shock has no effect on nonresidential investment, and a negative impact on real wages on impact. The right panel depicts impulse responses to an investment shock. In this case, nonresidential investment rises for three quarters following the shock, while the other responses are similar to those observed for a demand shock.
@@
Figure 10: The maximum $\hat{R}$ statistic for samples of different sizes from the elliptical slice sampler and NUTS. Horizontal dashed lines are located at 1.00 and 1.01. The values are calculated for the elements of $B$ for the NUTS, and the elements of $\Sigma$ and $Q$ for the elliptical slice sampler.
Figure 12: The minimum of bulk and tail ESS statistics for samples of different sizes of the elliptical slice sampler and NUTS. The values are calculated for elements of $B$ for the NUTS, and $\Sigma$ and $Q$ for the elliptical slice sampler.
@@
Figure 10 plots the maximum values of the $\hat{R}$ statistic for samples obtained using both samplers. The $\hat{R}$ is calculated for the elements of $\Sigma$ and $Q$ for the elliptical slice sampler, and for the elements of $B$ in the case of the NUTS. As in the previous application, we consider a threshold value of 1.01. As can be seen in the figure, $\hat{R}$ converges much faster when sampling using the NUTS. The elliptical slice sampler requires roughly 370 thousand saved draws to reach the threshold of 1.01, while the NUTS only needs 4 thousand total iterations. Hence, the NUTS takes approximately 5 minutes to obtain an acceptable sample, whereas the elliptical slice sampler requires roughly 46 minutes.
@@
The large number of iterations required by the elliptical slice sampler to converge suggests a high degree of autocorrelation also in this empirical application. Indeed, the chain generated by the elliptical slice sampler is highly autocorrelated, as Figure 11 shows. The autocorrelation function for the $(1,1)$ element of matrix $Q$ crosses zero only after approximately 700 lags, while the chain for the $(1,1)$ element of $B_{11}$ generated by the NUTS is uncorrelated beyond a few dozen lags.
@@
In this large SVAR, the efficiency per iteration of our method in terms of bulk and tail ESS for individual parameters is significantly better than that of the elliptical slice sampler, as indicated by Figure 12. After 100,000 iterations, the lowest bulk ESS for the parameters in the impact matrix $B$ is 25,700 using the NUTS. In contrast, the slice sampler achieves a minimum ESS of 520 for the elements of $\Sigma$ and $Q$ in 400 thousand iterations. In the case of tail ESS, the corresponding values are 35,650 for the NUTS and 1450 for the slice sampler.
As in Subsection 5.1, we also assess convergence using the minimum multivariate ESS of [vats2019] (the multivariate ESS is calculated for the parameters in columns 1–10 corresponding to the identified shocks of interest). To achieve a precision of 0.05 for the 95% confidence region, the NUTS requires roughly 7,000 iterations, whereas the elliptical slice sampler requires 126,000 iterations. These figures correspond to runtimes of 9 and 15,5 minutes, respectively.
@@
The time required to obtain 1,000 effective samples is 0.019 hours for the NUTS, compared to 0.047 hours for our implementation of the elliptical slice sampler. This is, however, slower than the runtime of less than one minute reported by [arias2025elliptical]. Our approach is faster than the algorithm of [chan2025], although their results are not directly comparable to ours, as they use the asymmetric conjugate prior of [chan2022], and only identify eight shocks instead of ten. They report a runtime of 140 minutes to obtain 1,000 draws, and as their draws are independent, this corresponds to an ESS of 1,000.
@@
To demonstrate that our approach can simultaneously incorporate both sign and zero restrictions, we augment the identification scheme with the restriction that labor productivity and utilization-adjusted TFP react on impact only to the technology and oil price shocks (assumed to have a positive impact effect). As discussed in Section 4.2, introducing these zero restrictions does not require importance sampling to ensure that the sample is from the correct distribution. Instead, the restricted values are simply fixed in the density function and omitted from the parameter vector. Hence, estimation time is largely unchanged by the addition of these restrictions.
In general, the impulse responses depicted in Figure 8 remain intact. However, the responses to the labor supply shock shown in Figure 13 change substantially when the zero restrictions are imposed. Under the baseline identification (left panel), labor productivity responds positively on impact and over the following three quarters, while the impact response of TFP is centered around zero but there is a lot of uncertainty about this estimate. In contrast, under the zero restrictions, the 68% posterior credible intervals for both variables include zero at all horizons.
Estimation of the model with additional exclusion restrictions remains efficient: the algorithm reaches a minimum of 19,700 and 31,900 bulk and tail ESS, respectively, while the maximum $\hat{R}$ crosses the 1.01 threshold after 4 thousand iterations, or 5 minutes. In terms of multivariate ESS, the sampling efficiency comes at 0.019 hours for 1,000 effective draws (as before, the effective draws are calculated for the impulse response parameters in columns 1–10 corresponding to the identified shocks of interest). These numbers are close to the case with no zero restrictions incorporated in the model.
@@ 6 Conclusion
Tight identifying inequality restrictions may render conventional accept-reject algorithms inefficient and in some cases infeasible for inference in SVAR models. In large-scale models that are becoming increasingly popular in empirical research, the high-dimensional parameter space and the large number of imposed restrictions can give rise to problems, even when individual restrictions are relatively weak, as their cumulative effect may still be substantial. In this paper, we propose an approach based on imposing inequality restrictions through a reparameterization of the structural model by continuously differentiable mappings. This approach accommodates various kinds of inequality restrictions, including shape and ranking restrictions on impulse responses at multiple horizons, as well as bounds on elasticities. Moreover, in contrast to the approaches of [arias2025elliptical, chan2025, read2025], our framework allows for a straightforward implementation of exclusion restrictions, with no significant increase in the speed of the sampler. As an additional advantage, the prior can be specified very flexibly, as our approach does not rely on direct sampling procedures, albeit this usually entails an increase in computation time.
@@ Appendix A Proof of Proposition 1
The construction of the Jacobian determinant turns out to be rather straightforward. Due to the relationship $B=\Psi_{0}$ and the fact that the formula in (A.1) is recursive, it follows that the Jacobian is a block-triangular matrix. Thus its determinant is the product of the determinants of the diagonal blocks. The first two diagonal blocks of dimensions $N\times N$ and $N^{2}\times N^{2}$ correspond to the identify maps on $\tilde{c}$ and $B$, respectively, and are thus identity matrices. The same is true for the $N^{2}\times N^{2}$ blocks in positions $k+1,\dots,p$, as they correspond to identity maps on matrices $\tilde{A}_{i}$. The remaining blocks on the diagonal correspond to the quantities
@@ Appendix C Estimation Algorithm
To implement an HMC sampler in practice, we use the Stan software package [stanreference], which implements an HMC variant known as the No-U-Turn sampler (NUTS), developed by [hoffman2014]. This algorithm tunes the number of steps of the leapfrog integrator dynamically by doubling the length of the path of the leapfrog integrator until the path starts to make a U-turn, at which point the expansion is stopped. Furthermore Stan implements procedures for adapting both $\epsilon$ and $M$ during the warmup phase of the chain, thus eliminating the need to hand-tune the values. $M$ can be specified to be either diagonal or a dense matrix, and the theoretically optimal value for $M$ equals the inverse of the covariance matrix of the target distribution [betancourt2018, Section 4.3]. For numerical reasons, practical implementations of HMC tend to work with the inverse metric $M^{-1}$ directly.
@@ Appendix D Non-centered parametrization
This parametrization alleviates the challenging geometry of the original posterior, as the part corresponding to the VAR matrices is standard normal, and thus the challenging funnel geometry is avoided completely. For the HMC algorithm, it is thus possible to run the leapfrog integrator with a larger stepsize and thus for fewer steps. Furthermore the evaluation of the density function of the standard normal distribution is much faster than the evaluation of a general multivariate normal distribution, which saves a lot on computation.
Note that $B$ and $Z$ are independent because $\Sigma$ and the auxiliary parameter $Z$ are independent and the structural impact matrix $B$ is a function of $\Sigma$ but not $Z$. For HMC this means that the inverse metric $M^{-1}$ (see Appendix C) is block-diagonal, as there are no covariances between $B$ and $Z$ to be taken into account. In our experience, this makes a diagonal inverse metric a viable choice when a non-centered parameterziation is used. Especially in a large model, estimating the inverse metric $M^{-1}$ is usually prohibitively time consuming, and sampling efficiency can be reduced if its estimate is far from the optimal value of the covariance matrix of the target posterior distribution [betancourt2018, Section 4.3]. Therefore, in the application of Section 5.2 involving the 35-variable model, we use a diagonal inverse metric, while in the application to the oil market in Section 5.1 a dense inverse metric is used because a non-centered parameterization is not feasible.

### 21 May 2026 v2

@@ 5.1 Small Oil Market Model
As an example of a large SVAR model, we consider a 35-variable quarterly model of the U.S. economy with five lags and a constant. This application extends the work of [chan2025], which in turn builds on [crump2021]. Following [arias2025elliptical], we consider the sample period from 1977:Q4 to 2019:Q4.111111We thank Christian Matthes for sharing their replication files. The variables are listed in Table 3, which also gives the identifying restrictions. Following [arias2025elliptical], both sign and ranking restrictions are imposed to identify ten structural shocks: demand, investment, financial, monetary, government spending, technology, labor supply, wage bargaining, oil price, and consumer sentiment shocks.
Following [arias2025elliptical], we place a Minnesota prior on the reduced-form parameters. [^12] We thank Jonas Arias for sharing their prior with us. As the model contains only contemporaneous restrictions, we estimate it using the non-centered specification (discussed brifely in Section 4.3), where instead of sampling the VAR coefficient matrices, we sample an auxiliary standard normal vector and transform it into the desired parameters $\tilde{A}$. Further details are provided in the Appendix D.
@@
Figure 7 depicts the impulse response functions of the real GDP, Fed funds rate, unemployment rate, PCE index, nonresidential investment, and real wage to a one standard deviation positive demand and investment shocks. The impulse responses are visually identical to those obtained using the elliptical slice sampler of [arias2025elliptical] under the same prior distribution and identifying restrictions, which suggests that our sampling algorithm correctly targets the same posterior distribution. In particular, the left panel of Figure 7 shows that the demand shock causes a temporary increase in the real GDP, prices, and the federal funds rate, but reduces unemployment. The demand shock has no effect on nonresidential investment nor real wages. The right panel depicts impulse responses to the investment shock, which leads to a rise in nonresidential investment rises for three quarters, while the other responses are similar to those observed for the demand shock.
@@
Figure 9: The maximum $\hat{R}$ statistic for samples of different sizes from the elliptical slice sampler and NUTS. Horizontal dashed lines are located at 1.00 and 1.01. The values are calculated for the elements of $B$ for the NUTS, and the elements of $\Sigma$ and $Q$ for the elliptical slice sampler. The vertical axis is restricted to $(1.0,1.1)$ for clarity.
Figure 10: Autocorrelation plots of the sample of $B_{11}$ obtained using the NUTS (top panel), and $Q_{11}$ obtained using the elliptical slice sampler (bottom panel).
@@
Figure 9 plots the maximum values of the $\hat{R}$ statistic for samples obtained using both samplers. The $\hat{R}$ is calculated for the elements of $\Sigma$ and $Q$ for the elliptical slice sampler, and for the elements of $B$ in the case of the NUTS. As in the previous application, we consider a threshold value of 1.01. As can be seen, $\hat{R}$ converges much faster when sampling using the NUTS. The elliptical slice sampler requires roughly 200 thousand saved draws to reach the threshold of 1.01, while the NUTS only needs five thousand total iterations. Hence, the NUTS takes approximately six minutes to obtain an acceptable sample, whereas the elliptical slice sampler requires approximately 24 minutes.
@@
The large number of iterations required by the elliptical slice sampler to converge suggests a high degree of autocorrelation also in this empirical application. Indeed, the chain generated by the elliptical slice sampler is highly autocorrelated, as Figure 10 shows. The autocorrelation coefficients for the $(1,1)$ element of the matrix $Q$ are positive for the first thousand lags, while the chain for the $(1,1)$ element of $B_{11}$ generated by the NUTS is uncorrelated beyond a few dozen lags.
Table 4: Minimum ESS measures per 1000 iterations. The values are calculated for elements of $B$ for the NUTS, and $\Sigma$ and $Q$ for the elliptical slice sampler.
ESS type
Sampler Bulk Tail
NUTS 217.21 341.21
Slice Sampler 2.24 5.11
@@
In this large SVAR, the efficiency per iteration of our method in terms of bulk and tail ESS for individual parameters is significantly better than that of the elliptical slice sampler, as indicated by Table 4. The minimum bulk ESS generated per 1000 iterations for the elements of $B$ is 217.21 when using the NUTS. In contrast, the minimum bulk ESS generated by the slice sampler for the elements of $\Sigma$ and $Q$ is 2.24. In the case of tail ESS, the corresponding values are 341.21 for the NUTS and 5.11 for the slice sampler.
@@
As in Section 5.1, we also assess convergence using the minimum multivariate ESS of [vats2019] (the multivariate ESS is calculated for the parameters in columns 1–10 corresponding to the identified shocks of interest). To achieve a precision of 0.05 for the 95% confidence region, the NUTS requires roughly 7,000 iterations, whereas the elliptical slice sampler requires 124,000 iterations. These figures correspond to runtimes of 8 and 15 minutes, respectively. The time required to obtain 1,000 effective samples is 1.24 minutes for the NUTS, compared to 2.64 minutes for our implementation of the elliptical slice sampler. Our approach is faster than the algorithm of [chan2025], although their results are not directly comparable to ours, as they use the asymmetric conjugate prior of [chan2022], and only identify eight shocks instead of ten. They report a runtime of 140 minutes to obtain 1,000 draws, and as their draws are independent, this corresponds to an ESS of 1,000.
@@
To demonstrate that our approach can simultaneously incorporate both sign and zero restrictions, we augment the identification scheme with the restriction that labor productivity and utilization-adjusted TFP react on impact only to the technology and oil price shocks (assumed to have a positive impact effect). As discussed in Section 4.2, introducing these zero restrictions does not require importance sampling to ensure that the sample is from the correct distribution. Instead, the restricted values are simply fixed in the density function and omitted from the parameter vector. Hence, estimation time is largely unchanged by the addition of these restrictions. The bulk and tail effective sample sizes per 1,000 iterations generated by the algorithm are 127.49 and 271.75, respectively, while the maximum $\hat{R}$ crosses the 1.01 threshold after 5 thousand iterations, or 7 minutes. In terms of the multivariate ESS, the sampling efficiency comes at 1.72 minutes for 1,000 effective draws (as before, the effective draws are calculated for the impulse response parameters in columns 1–10 corresponding to the identified shocks of interest). Thus, sampling is less efficient than in the case of having no zero restrictions, but the convergence time in terms of the $\hat{R}$ is essentially identical.
In general, the impulse responses depicted in Figure 7 remain intact. However, the responses to the labor supply shock shown in Figure 11 change substantially when the zero restrictions are imposed. Under the baseline identification (left panel), the response of labor productivity has a substantial amount of posterior mass above zero, while the impact response of TFP is centered around zero but there is a lot of uncertainty about this estimate. In contrast, under the zero restrictions, the 68% posterior credible intervals are much narrower at short horizons.
@@ 6 Conclusion
Tight identifying inequality restrictions may render conventional accept-reject algorithms inefficient and in some cases infeasible for inference in SVAR models. In large-scale models that are becoming increasingly popular in empirical research, the high-dimensional parameter space and the large number of imposed restrictions can give rise to problems, even when individual restrictions are relatively weak, as their cumulative effect may still be substantial. In this paper, we propose an approach based on imposing inequality restrictions through a reparameterization of the structural model by continuously differentiable mappings. This approach accommodates various kinds of inequality restrictions, including shape and ranking restrictions on impulse responses at multiple horizons, as well as bounds on elasticities. Moreover, in contrast to the approaches of [arias2025elliptical, chan2025, read2025], our framework allows for a straightforward implementation of exclusion restrictions, with no significant increase in the speed of the sampler. As an additional advantage, the prior can be specified in a very flexible manner, as our approach does not rely on direct sampling procedures, albeit this usually entails an increase in computation time.
@@ Appendix A Proof of Proposition 1
The construction of the Jacobian determinant turns out to be rather straightforward. Due to the relationship $B=\Psi_{0}$ and the fact that the formula in (A.1) is recursive, it follows that the Jacobian is a block-triangular matrix. Thus its determinant is the product of the determinants of the diagonal blocks. The first two diagonal blocks of dimensions $N\times N$ and $N^{2}\times N^{2}$ correspond to the identity maps on $\tilde{c}$ and $B$, respectively, and are thus identity matrices. The same is true for the $N^{2}\times N^{2}$ blocks in positions $k+1,\dots,p$, as they correspond to identity maps on matrices $\tilde{A}_{i}$. The remaining blocks on the diagonal correspond to the quantities
@@ Appendix C Estimation Algorithm
To implement an HMC sampler in practice, in this paper, we use the Stan software package [stanreference], which implements an HMC variant known as the No-U-Turn sampler (NUTS), developed by [hoffman2014]. This algorithm tunes the number of steps of the leapfrog integrator dynamically by doubling the length of the path of the leapfrog integrator until the path starts to make a U-turn, at which point the expansion is stopped. Furthermore Stan implements procedures for adapting both $\epsilon$ and $M$ during the warmup phase of the chain, thus eliminating the need to hand-tune the values. $M$ can be specified to be either diagonal or a dense matrix, and the theoretically optimal value for $M$ equals the inverse of the covariance matrix of the target distribution [betancourt2018, Section 4.3]. For numerical reasons, practical implementations of HMC tend to work with the inverse metric $M^{-1}$ directly.
@@ Appendix D Non-centered parametrization
This parametrization alleviates the challenging geometry of the original posterior, as the part corresponding to the VAR matrices is standard normal, and thus the challenging funnel geometry is avoided completely. For an HMC algorithm, it is thus possible to run the leapfrog integrator with a larger stepsize and thus for fewer steps. Furthermore, the evaluation of the density function of the standard normal distribution is much faster than the evaluation of a general multivariate normal distribution, which saves a lot on computation.
Note that $B$ and $Z$ are independent because $\Sigma$ and the auxiliary parameter $Z$ are independent and the structural impact matrix $B$ is a function of $\Sigma$ but not $Z$. For HMC this means that the inverse metric $M^{-1}$ (see Appendix C) is block-diagonal, as there are no covariances between $B$ and $Z$ to be taken into account. In our experience, this makes a diagonal inverse metric a viable choice when a non-centered parameterization is used. Especially in a large model, estimating the inverse metric $M^{-1}$ is usually prohibitively time consuming, and sampling efficiency can be reduced if its estimate is far from the optimal value of the covariance matrix of the target posterior distribution [betancourt2018, Section 4.3]. Therefore, in the application of Section 5.2 involving the 35-variable model, we use a diagonal inverse metric. In the oil market application in Section 5.1 we use a dense inverse metric because the distribution of $(\Psi_{1}\dots,\Psi_{12},\tilde{A}_{13},\dots,\tilde{A}_{24})$ is not a Gaussian, which precludes a non-centered parameterization.